Find an equation of the line that passes through the two given points. Write the equation in slope-intercept form, if possible
passes through (1, 8) and (−2, −1)

Answer by ewatrrr(24785)

(1, 8) and
(−2, −1) m = 9/3 = 3
...
y +1 = 3(x+2)
y = 3x + 5